# Sharp asymptotics for $N$-point correlation functions of coalescing heavy-tailed random walks

### v1

@@ 2.3 First hitting time
${\mathbb{P}}(\tau_{x}>t,X^{x}_{t}=y)\leq\frac{Ca(x)}{B_{t}\hskip 1.0ptl(t)}\quad\mbox{and}\quad\sum_{x}p(x){\mathbb{P}}(\tau_{x}>t,X^{x}_{t}=y)\leq\frac{C}{B_{t}\hskip 1.0ptl(t)}.$ (2.34)
Moreover, for any $\kappa<\alpha/(1-\alpha^{2})$, ((((if $\alpha=1$ then $\kappa<\infty$)))),
$\sum_{x}p(x)\Big{|}{\mathbb{P}}(\tau_{x}>t,X^{x}_{t}=y)-{\mathbb{P}}(\tau_{x}>t){\mathbb{P}}(X^{0}_{t}=0)\Big{|}\leq\frac{C}{B_{t}l(t)\ell(t)^{\kappa}}+\frac{C|y|}{B_{t}^{2}}.$ (2.35)
Proof For the bound in (2.34), we decompose at time $t/2$, By the Markov property,
@@
For the estimate in (2.35), note that if $|x|>B_{t}^{2}$, then for any $\varepsilon>0$,
$\sum_{|z|>B_{t}^{2}}p(x){\mathbb{P}}_{x}(\tau_{x}>t,X^{x}_{t}=y)\leq\sum_{|z|>B_{t}^{2}}p(x)\leq\frac{CL(B_{t}^{2})}{B_{t}^{2}}\leq\frac{C}{B_{t}^{2-\varepsilon}}.$ (2.36)
It remains to treat $|x|\leq B_{t}^{2}$. Let $H_{1}=H_{1}(t),H_{2}=H_{2}(t)$ be such that $t\gg H_{1}\gg 1$ and $H_{2}\gg 1$ At time $s=t/H_{1}$, we partition the space into two regions $D_{1}:=\{z:|z-x|>B_{s}H_{2}\}$ and $D_{2}:=\{z:|z-x|\leq B_{s}H_{2}\}$. By the uniform bound and the large deviation,
${\mathbb{P}}(\tau_{x}>t,X_{s}\in D_{1},X_{t}=y)\leq\sum_{z\in D_{1}}{\mathbb{P}}(X^{x}_{s}=z){\mathbb{P}}(X^{z}_{t-s}=y)\leq\frac{C}{B_{t}H_{2}^{\alpha-\varepsilon}}.$ (2.37)
@@
$\begin{array}[]{l}\displaystyle{\mathbb{P}}(\tau_{x}>t,X_{s}\in D_{2},X^{x}_{t}=y)\leq\sum_{z\in D_{2}}{\mathbb{P}}_{x}(\tau_{x}>t,X_{s}=z){\mathbb{P}}_{z}(X_{t-s}=y)\\[5.0pt] \displaystyle\qquad\leq\sum_{z\in{\mathbb{Z}}\setminus D_{1}}{\mathbb{P}}_{x}(\tau_{x}>t,X_{s}=z)\Big{(}{\mathbb{P}}_{0}(X_{t-s}=0)+\frac{C(|y|+|x|+B_{s}H_{2})}{B_{t}^{2}}\Big{)}\\[5.0pt] \displaystyle\qquad\leq\Big{(}{\mathbb{P}}_{x}(\tau_{x}>t)+\frac{C}{H_{2}^{\alpha-\varepsilon}}\Big{)}\Big{(}{\mathbb{P}}_{0}(X_{t}=0)+\frac{C}{B_{t}H_{1}^{\frac{1}{1+\alpha}-\varepsilon}}+\frac{C(|y|+|x|+B_{s}H_{2})}{B_{t}^{2}}\Big{)}\end{array}$ (2.38)
@@
$B_{t}^{-2}\sum_{|x|\leq B_{t}^{2}}|x|p(x)\leq B_{t}^{-2}\sum_{x}B_{t}^{\varepsilon}\hskip 1.0pt|x|^{1-\varepsilon/2}p(x)\leq CB_{t}^{-2+\varepsilon}.$ (2.39)
For the other error terms, fix any $\tilde{\kappa}<\alpha/(1-\alpha)$, and set $H_{1}=l(t)^{3/\alpha}\ell(t)^{\tilde{\kappa}}$ (so that $H_{1}\ll t$) and $H_{2}=H_{1}^{\frac{1}{\alpha(1+\alpha)}}$. Then, using Lemma 2.9, direct calculations and basic estimates yield that
$\sum_{x}p(x)\Big{|}{\mathbb{P}}_{x}(\tau_{x}>t,X_{t}=y)-{\mathbb{P}}_{x}(\tau_{x}>t){\mathbb{P}}_{0}(X_{t}=0)\Big{|}\leq\frac{C}{B_{t}l(t)\ell(t)^{\kappa}}+\frac{C|y|}{B_{t}^{2}},$ (2.40)
where $\kappa$ can be made arbitrary close to $\tilde{\kappa}/(1+\alpha)$ (and hence to $\alpha/(1-\alpha^{2})$) as $\varepsilon\to 0$.
@@ 3 First collision time of random walks
Recall from Theorem 1.4 that $({\mathbf{X}}_{t})_{t\geq 0}=(X^{1}_{t},\ldots,X^{N}_{t})_{t\geq 0}$ represents $N$ independent random walks with initial state ${\mathbf{X}}_{0}={\bf x}\in{\mathbb{Z}}^{N}$ and jump kernel $p(\cdot)$ in (1.4). Recall from (1.11) that $\tau_{{\bf x}}$ is the first collision time of ${\mathbf{X}}$. We show Theorem 1.4 for asymptotic probability of the non-collision event $\{\tau_{{\bf x}}>t\}$ in Subsection 3.1, and moreover show that this event is asymptotically independent of the transition event $\{X_{t}={\bf y}\}$ in Subsection 3.2.
@@
As outlined in Subsection 1.2, our goal is to derive an ODE to characterize the asymptotic behavior of $p_{\rm NC}(t)={\mathbb{P}}(\tau_{{\bf x}}>t)$. The generator calculation yields (1.18), i.e.,
$-\frac{\mathrm{d}}{\mathrm{d}t}p_{\rm NC}(t)=\sum_{i<j}\sum_{y_{ij}\in{\mathbb{Z}}}2p(y_{ij}){\mathbb{P}}(\tau_{{\bf x}}>t,\Delta^{ij}_{t}=y_{ij}),$ (3.1)
where $\Delta^{ij}_{t}=X^{i}_{t}-X^{j}_{t}$. Our first step is to approximate the r.h.s. using the lemma below. Recall the function class ${\cal B}$ such that $B_{t}\in{\cal B}$ implies $\ell(t)^{K}\geq l(t)$ for some $K>0$.
@@
Assume that $B_{t}$ belongs to the class ${\cal B}$. Let $\kappa>0$. Let $s=t/\ell(t)^{\theta}$ for some $\theta>K+\kappa+1$ if $\alpha=1$ or $\theta<\frac{1}{1-\alpha}$ if $\alpha<1$. Then, there exist $c=c({\bf x})$ and $\delta>0$ such that for any $\varepsilon>0$,
$\sum_{y_{ij}}p(y_{ij})\big{|}{\mathbb{P}}(\tau_{\bf x}>t,\Delta^{ij}_{t}=y_{ij})-\frac{p_{\rm NC}(s)}{B_{2t}\hskip 1.0ptl(2t)}\big{|}\leq\left\{\begin{array}[]{ll}\displaystyle\frac{c}{B_{t}\hskip 1.0ptl(t)}\Big{(}\frac{p_{\rm NC}(s)}{\ell(t)^{1-\varepsilon}}+\frac{1}{\ell(t)^{\kappa}}\Big{)},&\alpha=1;\\[10.0pt] \displaystyle\frac{c}{B_{t}\hskip 1.0ptl(t)t^{\delta}},&\alpha\in(0,1).\end{array}\right.$ (3.2)
@@
$\begin{array}[]{l}\displaystyle\quad\sum_{y_{ij}}p(y_{ij})\big{|}{\mathbb{P}}(\tau_{\bf x}>t,\Delta^{ij}_{t}=y_{ij})-{\mathbb{P}}(\tau_{\bf x}>s,\tau_{ij}>t,\Delta^{ij}_{t}=y_{ij})\big{|}\\[5.0pt] \displaystyle\leq\sum_{\{k,l\}\neq\{i,j\}}\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}(\tau_{\bf x}>s,\tau_{ij}>t,\Delta^{ij}_{t}=y_{ij},s<\tau_{kl}\leq t).\end{array}$ (3.3)
We bound the r.h.s. of (3.3) by decomposing it into three cases as: (a) $\{\tau_{kl}>s+u\}$, (b)$\{\tau_{kl}\leq s+u,\hskip 2.0pt|\Delta^{kl}_{s}|>B_{u}/\ell(u)^{\theta_{1}}\}$, (c) $\{\tau_{kl}\leq s+u,\hskip 2.0pt|\Delta^{kl}_{s}|\leq B_{u}/\ell(u)^{\theta_{1}}\}$. Here, we take $u=t/2$ if $\alpha=1$ and $u=t/\ell(t)^{\theta_{2}}$ otherwise, where $\theta_{1}>0,\theta_{2}\in(0,\frac{1}{1-\alpha})$ will be chosen later.
On the intersection with (a), we decompose at the collision time $\tau_{kl}.$ Let $v:=\tau_{kl}-s$. For ${\mathbf{X}}_{s}={\bf z}\in{\mathbb{Z}}^{N}$, let $({\mathbf{X}}_{t}^{z})_{t\geq 0}:=({\mathbf{X}}_{t+s})_{t\geq 0}$ denote the continuation of the random walks starting from time $s$. We also use ${\mathbb{P}}_{\bf z}$ to denote the law of random walks with a initial state ${\bf z}$. Conditioning on ${\mathbf{X}}_{s}={\bf z}\in{\mathbb{Z}}^{N}$, the random walks first evolve to ${\mathbf{X}}^{{\bf z}}_{v}=\textbf{w}$, with $\Delta_{v}^{kl}=w_{kl}=w_{k}-w_{l}$. Then at time $v$, $\Delta^{kl}$ jumps to 0 at rate $2p(w_{kl})$. Over the remaining time $t-s-v$, $\Delta^{ij}$ moves from $w_{ij}$ to $y_{ij}$ without hitting 0. Thus, the contribution from (a) in the r.h.s. of (3.3) is bounded above by
$\begin{array}[]{l}\displaystyle\sum_{\{k,l\}\neq\{i,j\}}\int_{u}^{t-s}\sum_{{\bf z}}{\mathbb{P}}(\tau_{{\bf x}}>s,{\mathbf{X}}_{s}={\bf z})\sum_{\textbf{w}}{\mathbb{P}}_{{\bf z}}(\tau_{kl}>v,{\mathbf{X}}_{v}=\textbf{w})\\[5.0pt] \displaystyle\qquad\qquad\qquad\qquad\times 2p(w_{kl})\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{w_{ij}}(\tau_{ij}>t-s-v,\Delta^{ij}_{t-s-v}=y_{ij})\mathrm{d}v.\end{array}$ (3.4)
@@
${\mathbb{P}}_{\bf z}(\tau_{kl}>v,{\mathbf{X}}^{\{i\}}_{v}=\textbf{w}^{\{i\}},X^{i}_{v}=w_{j}+w_{ij})\leq\frac{C}{B_{v}}{\mathbb{P}}_{\bf z}(\tau_{kl}>v,{\mathbf{X}}^{\{i\}}_{v}=\textbf{w}^{\{i\}}).$ (3.5)
Now for the event in the last sum in (3.4), it is equivalent to that the time-reversed random walk $\widehat{\Delta}^{ij}$ moves from $y_{ij}$ to $w_{ij}$ during a time period $t-s-v=:r$ and $\hat{\tau}_{ij}>r$. Thus,
$\sum_{w_{ij}}\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{w_{ij}}(\tau_{ij}>r,\Delta^{ij}_{r}=y_{ij})\displaystyle=\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{y_{ij}}(\hat{\tau}_{ij}>r)\leq\frac{C}{l(r)}.$ (3.6)
Inserting (3.5)-(3.6) into (3.4) and using a similar time-reversal argument, we obtain an upper for the middle sum in (3.4) over $\textbf{w}^{\{i\}}$ as
$\sum_{\textbf{w}^{\{i\}}}p(w_{kl}){\mathbb{P}}_{{\bf z}}(\tau_{kl}>v,{\mathbf{X}}^{\{i\}}_{v}=\textbf{w}^{\{i\}})=\sum_{w_{kl}}p(w_{kl}){\mathbb{P}}_{w_{kl}}(\hat{\tau}_{kl}>v,\widehat{\Delta}^{kl}_{v}=z_{kl})\leq\frac{C}{B_{v}l(v)},$ (3.7)
where the inequality is due to Lemma 2.12. Consequently, (3.4) has an upper bound
$C_{N}\int_{u}^{t-s}\frac{1}{B^{2}_{v}\hskip 1.0ptl(v)\hskip 1.0ptl(t-s-v)}\sum_{{\bf z}}{\mathbb{P}}(\tau_{{\bf x}}>s,{\mathbf{X}}={\bf z})\mathrm{d}v=C_{N}\int_{u}^{t-s}\frac{p_{\rm NC}(s)}{B^{2}_{v}\hskip 1.0ptl(v)\hskip 1.0ptl(t-s-v)}\mathrm{d}v.$ (3.8)
Note that $t\gg s$ and $t/2\geq u$. Split the integral domain into $(u,2t/3)$ where $l(t-s-v)\asymp l(t)$, and $(2t/3,t-s)$ where $B_{v}^{2}\hskip 1.0ptl(v)\asymp B_{t}^{2}\hskip 1.0ptl(t)$. Applying the Potter bound and Karamata’s theorem for the integrals then yields that for any $\varepsilon>0$, (3.4) is bounded above by
$C_{N}\frac{t}{B^{2}_{t}\hskip 1.0ptl(t)^{2}}\hskip 1.0pt(t/u)^{\frac{2}{\alpha}-1+\varepsilon}p_{\rm NC}(s)=\frac{C_{N}(t/u)^{\frac{2}{\alpha}-1+\varepsilon}}{B_{t}\hskip 1.0ptl(t)\ell(t)}\hskip 1.0ptp_{\rm NC}(s).$ (3.9)
On (b), during $(s+u,t]$, $\Delta^{ij}$ moves from some $w_{ij}\in{\mathbb{Z}}$ to $y_{ij}$ without hitting 0. The time-reversal argument and Lemma 2.13 imply that uniformly in $w_{ij}$,
$\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{w_{ij}}(\Delta^{ij}_{\tilde{r}}=y_{ij},\tau>\tilde{r})=\displaystyle\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{y_{ij}}(\widehat{\Delta}^{ij}_{\tilde{r}}=w,\hat{\tau}>\tilde{r})\leq\frac{C}{B_{t}\hskip 1.0ptl(t)},$ (3.10)
@@
${\mathbb{P}}_{z_{kl}}(\tau_{kl}\leq u)\leq C\ell(u)^{-1+(1-\alpha)\theta_{1}+\varepsilon}.$ (3.11)
Applying the Markov property and using above bounds uniform in $w_{ij}$ and in $|z_{kl}|>B_{u}/\ell(u)^{\theta_{1}}$, we deduce that the contribution from (b) in the r.h.s. of (3.3) is bounded above by
$\sum_{\{k,l\}\neq\{i,j\}}\sum_{|z_{kl}|>B_{u}/\ell(u)^{\theta_{1}}}{\mathbb{P}}(\tau_{\bf x}>s,\Delta_{s}^{kl}=z_{kl})\frac{C\ell(u)^{(1-\alpha)\theta_{1}+\varepsilon}}{B_{t}\hskip 1.0ptl(t)\ell(u)}\leq\frac{C_{N}\ell(u)^{(1-\alpha)\theta_{1}+\varepsilon}}{B_{t}\hskip 1.0ptl(t)\ell(u)}p_{\rm NC}(s).$ (3.12)
On (c), use the same time-reversal argument as for (3.10) for the interval $(s,t]$, and on $[0,s]$ simply apply the uniform transition probability for $\Delta^{kl}$. This yields the upper bound
$\begin{array}[]{l}\displaystyle\sum_{\{k,l\}\neq\{i,j\}}\sum_{z_{ij}}{\mathbb{P}}(|\Delta_{s}^{kl}|\leq B_{u}/\ell(u)^{\theta_{1}},\Delta_{s}^{ij}=z_{ij})\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{y_{ij}}(\widehat{\Delta}^{ij}_{t-s}=z_{ij},\hat{\tau}_{ij}>t-s)\\[5.0pt] \hskip 56.9055pt\displaystyle\leq\frac{C_{N}B_{u}}{B_{s}\ell(s)^{\theta_{1}}}\hskip 1.0pt\frac{1}{B_{t}l(t)}.\end{array}$ (3.13)
Combining the three bounds from (3.9), (3.12) and (3.13), we deduce that
@@
For $\alpha=1$, recall $u=t/2$, and take $\theta_{1}=\kappa+2\theta$ so that $\frac{1}{B_{s}\ell(s)^{\theta_{1}}}\leq\frac{(t/s)^{1+\varepsilon}}{B_{t}\ell(t)^{\theta_{1}}}\leq\frac{1}{B_{t}\ell(t)^{\kappa}}$ by the Potter bound. Inserting this into (3.14), basic estimates then imply the first line in the r.h.s. of (3.2). For $\alpha\in(0,2)$ and $u=t/\ell(t)^{\theta_{2}}$, recall that $\ell$ is $(1/\alpha-1)$-regularly varying. Bounding $p_{\rm NC}(s)\leq 1$ and taking suitable $\theta_{1},\theta_{2}$, it is not hard to see that the second line in the r.h.s. of (3.2) holds for some $\delta>0$.

### v2

@@ 2.3 First hitting time
${\mathbb{P}}(\tau_{x}>t,X^{x}_{t}=y)\leq\frac{Ca(x)}{B_{t}\hskip 1.0ptl(t)}\quad\mbox{and}\quad\sum_{x}p(x){\mathbb{P}}(\tau_{x}>t,X^{x}_{t}=y)\leq\frac{C}{B_{t}\hskip 1.0ptl(t)}.$ (56)
Moreover, for any $\kappa<\alpha/(1-\alpha)$, ((if $\alpha=1$ then $\kappa<\infty$)),
$\sum_{x}p(x)\Big|{\mathbb{P}}(\tau_{x}>t,X^{x}_{t}=y)-{\mathbb{P}}(\tau_{x}>t){\mathbb{P}}(X^{0}_{t}=0)\Big|\leq\frac{C}{B_{t}l(t)\ell(t)^{\kappa}}+\frac{C|y|}{B_{t}^{2}}.$ (57)
Proof For the bound in (56), we decompose at time $t/2$, by the Markov property,
@@
For the estimate in (57), note that if $|x|>B_{t}^{2}$, then for any $\varepsilon>0$,
$\sum_{|z|>B_{t}^{2}}p(x){\mathbb{P}}(\tau_{x}>t,X^{x}_{t}=y)\leq\sum_{|z|>B_{t}^{2}}p(x)\leq\frac{CL(B_{t}^{2})}{B_{t}^{2}}\leq\frac{C}{B_{t}^{2-\varepsilon}}.$ (58)
It remains to treat $|x|\leq B_{t}^{2}$. Let $H_{1}=H_{1}(t),H_{2}=H_{2}(t)$ be such that $t\gg H_{1}\gg 1$ and $H_{2}\gg 1$ . At time $s=t/H_{1}$, we partition the space into two regions $D_{1}:=\{z:|z-x|>B_{s}H_{2}\}$ and $D_{2}:=\{z:|z-x|\leq B_{s}H_{2}\}$. By the uniform bound and the large deviation,
${\mathbb{P}}(\tau_{x}>t,X_{s}\in D_{1},X_{t}=y)\leq\sum_{z\in D_{1}}{\mathbb{P}}(X^{x}_{s}=z){\mathbb{P}}(X^{z}_{t-s}=y)\leq\frac{C}{B_{t}H_{2}^{\alpha-\varepsilon}}.$ (59)
@@
$\begin{array}[]{l}\displaystyle{\mathbb{P}}(\tau_{x}>t,X_{s}\in D_{2},X^{x}_{t}=y)\leq\sum_{z\in D_{2}}{\mathbb{P}}_{x}(\tau_{x}>t,X_{s}=z){\mathbb{P}}_{z}(X_{t-s}=y)\\[5.0pt] \displaystyle\qquad\leq\sum_{z\in{\mathbb{Z}}\setminus D_{1}}{\mathbb{P}}_{x}(\tau_{x}>t,X_{s}=z)\Big({\mathbb{P}}_{0}(X_{t-s}=0)+\frac{C(|y|+|x|+B_{s}H_{2})}{B_{t}^{2}}\Big)\\[5.0pt] \displaystyle\qquad\leq\Big({\mathbb{P}}_{x}(\tau_{x}>t)+\frac{C}{H_{2}^{\alpha-\varepsilon}}\Big)\Big({\mathbb{P}}_{0}(X_{t}=0)+\frac{C}{B_{t}H_{1}}+\frac{C(|y|+|x|+B_{s}H_{2})}{B_{t}^{2}}\Big)\end{array}$ (60)
@@
$B_{t}^{-2}\sum_{|x|\leq B_{t}^{2}}|x|p(x)\leq B_{t}^{-2}\sum_{x}B_{t}^{\varepsilon}\hskip 1.0pt|x|^{1-\varepsilon/2}p(x)\leq CB_{t}^{-2+\varepsilon}.$ (61)
For the other error terms, fix any $\kappa<\alpha/(1-\alpha)$, and set $H_{1}=l(t)^{3/\alpha}\ell(t)^{\kappa}$ (so that $H_{1}\ll t$) and $H_{2}=H_{1}^{1/\alpha}$. Then, using Lemma 2.9, direct calculations and basic estimates yield that
$\sum_{x}p(x)\Big|{\mathbb{P}}_{x}(\tau_{x}>t,X_{t}=y)-{\mathbb{P}}_{x}(\tau_{x}>t){\mathbb{P}}_{0}(X_{t}=0)\Big|\leq\frac{C}{B_{t}l(t)\ell(t)^{\kappa}}+\frac{C|y|}{B_{t}^{2}},$ (62)
where $\kappa$ can be made arbitrary close to $\alpha/(1-\alpha)$ as $\varepsilon\to 0$.
@@ 3 First collision time of random walks
Recall from Theorem 1.2 that $({\mathbf{X}}_{t})_{t\geq 0}=(X^{1}_{t},\ldots,X^{N}_{t})_{t\geq 0}$ represents $N$ independent random walks with initial state ${\mathbf{X}}_{0}={\bf x}\in{\mathbb{Z}}^{N}$ and jump kernel $p(\cdot)$ in (4). Recall from (10) that $\tau_{{\bf x}}$ is the first collision time of $\mathbf{X}$}. We show Theorem 1.2 for asymptotic probability of the non-collision event $\{\tau_{{\bf x}}>t\}$ in Subsection 3.1, and moreover show that this event is asymptotically independent of the transition event $\{X_{t}={\bf y}\}$ in Subsection 3.2.
@@
As outlined in Subsection 1.2, our goal is to derive an ODE to characterize the asymptotic behavior of $p_{\rm NC}(t)={\mathbb{P}}(\tau_{{\bf x}}>t)$. The generator calculation yields (19), i.e.,
$\frac{\mathrm{d}}{\mathrm{d}t}p_{\rm NC}(t)=\sum_{i<j}\sum_{y_{ij}\in{\mathbb{Z}}}2p(y_{ij}){\mathbb{P}}(\tau_{{\bf x}}>t,\Delta^{ij}_{t}=y_{ij}),$ (63)
[…]
@@
[…]
$\Big|\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}(\tau_{\bf x}>t,\Delta^{ij}_{t}=y_{ij})-\frac{p_{\rm NC}(s)}{B_{2t}\hskip 1.0ptl(2t)}\Big|\leq\left\{\begin{array}[]{ll}\displaystyle\frac{c}{B_{t}\hskip 1.0ptl(t)}\Big(\frac{p_{\rm NC}(s)}{\ell(t)^{1-\varepsilon}}+\frac{1}{\ell(t)^{\kappa}}\Big),&\alpha=1;\\[10.0pt] \displaystyle\frac{c}{B_{t}\hskip 1.0ptl(t)t^{\delta}},&\alpha\in(0,1).\end{array}\right$. (64)
@@
$\begin{array}[]{l}\displaystyle\quad\sum_{y_{ij}}p(y_{ij})\big|{\mathbb{P}}(\tau_{\bf x}>t,\Delta^{ij}_{t}=y_{ij})-{\mathbb{P}}(\tau_{\bf x}>s,\tau_{ij}>t,\Delta^{ij}_{t}=y_{ij})\big|\\[5.0pt] \displaystyle\leq\sum_{\{k,l\}\neq\{i,j\}}\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}(\tau_{\bf x}>s,\tau_{ij}>t,\Delta^{ij}_{t}=y_{ij},s<\tau_{kl}\leq t).\end{array}$ (65)
[…]
On the intersection with (a), we decompose at the collision time $\tau_{kl}.$ Let $v:=\tau_{kl}-s$. For ${\mathbf{X}}_{s}={\bf z}\in{\mathbb{Z}}^{N}$, let $({\mathbf{X}}_{t}^{z})_{t\geq 0}:=({\mathbf{X}}_{t+s})_{t\geq 0}$ denote the continuation of the random walks starting from time $s$. We also use ${\mathbb{P}}_{\bf z}$ to denote the law of random walks with a initial state $\bf z$}. Conditioning on ${\mathbf{X}}_{s}={\bf z}\in{\mathbb{Z}}^{N}$, the random walks first evolve to ${\mathbf{X}}^{{\bf z}}_{v}=\textbf{w}$, with $\Delta_{v}^{kl}=w_{kl}=w_{k}-w_{l}$. Then at time $v$, $\Delta^{kl}$ jumps to 0 at rate $2p(w_{kl})$. Over the remaining time tsvt-s-v, $\Delta^{ij}$ moves from $w_{ij}$ to $y_{ij}$ without hitting 0. Thus, the contribution from (a) in the r.h.s. of (65) is bounded above by
$\begin{array}[]{l}\displaystyle\sum_{\{k,l\}\neq\{i,j\}}\int_{u}^{t-s}\sum_{{\bf z}}{\mathbb{P}}(\tau_{{\bf x}}>s,{\mathbf{X}}_{s}={\bf z})\sum_{\textbf{w}}{\mathbb{P}}_{{\bf z}}(\tau_{kl}>v,{\mathbf{X}}_{v}=\textbf{w})\\[5.0pt] \displaystyle\qquad\qquad\qquad\qquad\times 2p(w_{kl})\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{w_{ij}}(\tau_{ij}>t-s-v,\Delta^{ij}_{t-s-v}=y_{ij})\mathrm{d}v.\end{array}$ (66)
@@
${\mathbb{P}}_{\bf z}(\tau_{kl}>v,{\mathbf{X}}^{\{i\}}_{v}=\textbf{w}^{\{i\}},X^{i}_{v}=w_{j}+w_{ij})\leq\frac{C}{B_{v}}{\mathbb{P}}_{\bf z}(\tau_{kl}>v,{\mathbf{X}}^{\{i\}}_{v}=\textbf{w}^{\{i\}}).$ (67)
Now for the event in the last sum in (66), it is equivalent to that the time-reversed random walk $\widehat{\Delta}^{ij}$ moves from $y_{ij}$ to $w_{ij}$ during a time period tsv=:rt-s-v=:r and $\hat{\tau}_{ij}>r$. Thus,
$\sum_{w_{ij}}\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{w_{ij}}(\tau_{ij}>r,\Delta^{ij}_{r}=y_{ij})\displaystyle=\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{y_{ij}}(\hat{\tau}_{ij}>r)\leq\frac{C}{l(r)}.$ (68)
Inserting (67)-(68) into (66) and using a similar time-reversal argument, we obtain an upper bound for the middle sum in (66) over $\textbf{w}^{\{i\}}$ as
$\sum_{\textbf{w}^{\{i\}}}p(w_{kl}){\mathbb{P}}_{{\bf z}}(\tau_{kl}>v,{\mathbf{X}}^{\{i\}}_{v}=\textbf{w}^{\{i\}})=\sum_{w_{kl}}p(w_{kl}){\mathbb{P}}_{w_{kl}}(\hat{\tau}_{kl}>v,\widehat{\Delta}^{kl}_{v}=z_{kl})\leq\frac{C}{B_{v}l(v)},$ (69)
where the inequality is due to Lemma 2.12. Consequently, (66) has an upper bound
$C_{N}\int_{u}^{t-s}\frac{1}{B^{2}_{v}\hskip 1.0ptl(v)\hskip 1.0ptl(t-s-v)}\sum_{{\bf z}}{\mathbb{P}}(\tau_{{\bf x}}>s,{\mathbf{X}}={\bf z})\mathrm{d}v=C_{N}\int_{u}^{t-s}\frac{p_{\rm NC}(s)}{B^{2}_{v}\hskip 1.0ptl(v)\hskip 1.0ptl(t-s-v)}\mathrm{d}v.$ (70)
Note that $t\gg s$ and $t/2\geq u$. Split the integral domain into $(u,2t/3)$ where $l(t-s-v)\asymp l(t)$, and (2t/3,ts)(2t/3,t-s) where $B_{v}^{2}\hskip 1.0ptl(v)\asymp B_{t}^{2}\hskip 1.0ptl(t)$. Applying the Potter bound and Karamata’s theorem for the integrals then yields that for any $\varepsilon>0$, (66) is bounded above by
$C_{N}\frac{t}{B^{2}_{t}\hskip 1.0ptl(t)^{2}}\hskip 1.0pt(t/u)^{\frac{2}{\alpha}-1+\varepsilon}p_{\rm NC}(s)=\frac{C_{N}(t/u)^{\frac{2}{\alpha}-1+\varepsilon}}{B_{t}\hskip 1.0ptl(t)\ell(t)}\hskip 1.0ptp_{\rm NC}(s).$ (71)
On (b), during $(s+u,t]$, $\Delta^{ij}$ moves from some $w_{ij}\in{\mathbb{Z}}$ to $y_{ij}$ without hitting 0. By Markov’s property the r.h.s. of (65) is
$\begin{array}[]{l}\displaystyle\sum_{\{k,l\}\neq\{i,j\}}\sum_{|z_{kl}|>B_{u}/\ell(u)^{\theta_{1}}}\sum_{w_{ij}}{\mathbb{P}}(\tau_{\bf x}>s,s<\tau_{kl}\leq s+u,\Delta^{kl}_{s}=z_{kl},\Delta^{ij}_{s+u}=w_{ij})\\[5.0pt] \hskip 85.35826pt\displaystyle\times\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{w_{ij}}(\tau_{ij}>t-s-u,\Delta^{ij}_{t-s-u}=y_{ij})\end{array}$ (72)
The time-reversal argument and Lemma 2.13 imply that uniformly in $w_{ij}$,
$\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{w_{ij}}(\Delta^{ij}_{\tilde{r}}=y_{ij},\tau>\tilde{r})=\displaystyle\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{y_{ij}}(\widehat{\Delta}^{ij}_{\tilde{r}}=w,\hat{\tau}>\tilde{r})\leq\frac{C}{B_{t}\hskip 1.0ptl(t)},$ (73)
@@
${\mathbb{P}}_{z_{kl}}(\tau_{kl}\leq u)\leq C\ell(u)^{-1+(1-\alpha)\theta_{1}+\varepsilon}.$ (74)
Applying the Markov property and using above bounds uniform in $w_{ij}$ and in $|z_{kl}|>B_{u}/\ell(u)^{\theta_{1}}$, we deduce that the contribution from (b) in the r.h.s. of (65) is bounded above by
$\sum_{\{k,l\}\neq\{i,j\}}\sum_{|z_{kl}|>B_{u}/\ell(u)^{\theta_{1}}}{\mathbb{P}}(\tau_{\bf x}>s,\Delta_{s}^{kl}=z_{kl})\frac{C\ell(u)^{(1-\alpha)\theta_{1}+\varepsilon}}{B_{t}\hskip 1.0ptl(t)\ell(u)}\leq\frac{C_{N}\ell(u)^{(1-\alpha)\theta_{1}+\varepsilon}}{B_{t}\hskip 1.0ptl(t)\ell(u)}p_{\rm NC}(s).$ (75)
On (c), use the same time-reversal argument as for (73) for the interval $(s,t]$, and on $[0,s]$ simply apply the uniform transition probability for $\Delta^{kl}$. This yields the upper bound
$\begin{array}[]{l}\displaystyle\sum_{\{k,l\}\neq\{i,j\}}\sum_{z_{ij}}{\mathbb{P}}(|\Delta_{s}^{kl}|\leq B_{u}/\ell(u)^{\theta_{1}},\Delta_{s}^{ij}=z_{ij})\sum_{y_{ij}}p(y_{ij}){\mathbb{P}}_{y_{ij}}(\widehat{\Delta}^{ij}_{t-s}=z_{ij},\hat{\tau}_{ij}>t-s)\\[5.0pt] \hskip 56.9055pt\displaystyle\leq\frac{C_{N}B_{u}}{B_{s}\ell(u)^{\theta_{1}}}\hskip 1.0pt\frac{1}{B_{t}l(t)}.\end{array}$ (76)
Combining the three bounds from (71), (75) and (76), we deduce that
@@
For $\alpha=1$, recall $u=t/2$, and take $\theta_{1}=\kappa+2\theta$ so that $\frac{1}{B_{s}\ell(s)^{\theta_{1}}}\leq\frac{(t/s)^{1+\varepsilon}}{B_{t}\ell(t)^{\theta_{1}}}\leq\frac{1}{B_{t}\ell(t)^{\kappa}}$ by the Potter bound. Inserting this into (77), basic estimates then imply the first line in the r.h.s. of (64). For $\alpha\in(0,1)$ and $u=t/\ell(t)^{\theta_{2}}$, recall that $\ell$ is $(1/\alpha-1)$-regularly varying. Bounding $p_{\rm NC}(s)\leq 1$ and taking suitable $\theta_{1},\theta_{2}$, it is not hard to see that the second line in the r.h.s. of (64) holds for some $\delta>0$.